OpenStudy (anonymous):

Would you use the change base formula for this equation? (solving logarithmic and exponential equations)

OpenStudy (anonymous):

\[\log(5)x=3\]

OpenStudy (amistre64):

you can if you want ... and in latex: log_5 gets you to the base 5

OpenStudy (amistre64):

\[log_5x=3\]

OpenStudy (anonymous):

so how would you solve it. I need help understanding

OpenStudy (amistre64):

well, a log is just an exponent of its base. so i would solve it by basing each side by 5 \[\large 5^{log_5x}=5^3\] \[x=5^3\]

OpenStudy (amistre64):

if we start with: x = 5^3 we can log each side: log(x) = log(5^3) we can pull out the exponent: log(x) = 3 log(5) then divide off the log(5): log(x)/log(5) = 3 and the left side is just change of base: log_5 (x) = 3 this is a better way to see the solution if you want to change of base it

OpenStudy (anonymous):

Ohhh okay thank you

OpenStudy (amistre64):

youre welcome